OpenStudy (anonymous):

Found this one in one of my old test papers, give it a try folks: If we have a prime no \( p \) and natural nos \( x \) and \( y \) such that \( p^x = y^4 + 4 \), then how many such ordered triplets \( (p, x, y) \) exist?

OpenStudy (dumbcow):

i have no idea :)

OpenStudy (katrinakaif):

Why is everyone looking through their old papers now? :P

OpenStudy (anonymous):

across question reminded me I have to work on mine too :P

OpenStudy (anonymous):

no is "no" and nos is numbers.

OpenStudy (anonymous):

I think it has only one solution which is \((5,1,1)\).

OpenStudy (anonymous):

@AnwarA: The thing is .. you have to prove it ;)

OpenStudy (anonymous):

I know! Workin' on it ;)

OpenStudy (anonymous):

Cool! :)

jhonyy9 (jhonyy9):

\[(2n+1)^{x}=y ^{4}+4\]

jhonyy9 (jhonyy9):

- what is your opinion(s) from this ? xln(2n+1)=4ln(y)+ln4 xln(2n+1)=4ln(4y) - so i think this is equal when x=4 and 2n+1=4y so y=(2n+1)/4 so but we know that (2n+1) not can divide by 4 never because every primes grater than 2 are odd

OpenStudy (mathmate):

ln(y^4+4) does not equal 4ln(y)+ln(4)